### v2

# RACS and SADL: Towards Robust SMR in the Wide-Area Network

Abstract
Widely deployed consensus protocols in the cloud are often leader-based and optimized for low latency under synchronous network conditions. However, cloud networks can experience disruptions such as network partitions, high-loss links, and configuration errors. These disruptions interfere with the operation of leader-based protocols, as their view change mechanisms interrupt the normal case replication and cause the system to stall.
This paper proposes RACS, a novel randomized consensus protocol that ensures robustness against adversarial network conditions. RACS achieves optimal one-round trip latency under synchronous network conditions while remaining resilient to adversarial network conditions. RACS follows a simple design inspired by Raft, the most widely used consensus protocol in the cloud, and therefore enables seamless integration with the existing cloud software stack— a goal no previous asynchronous protocol has successfully achieved.
Experiments with a prototype deployed on Amazon EC2 confirm that RACS achieves a throughput of 28k cmd/sec under adversarial cloud network conditions, whereas existing leader-based protocols such as Multi-Paxos and Raft provide less than 2.8k cmd/sec. Under synchronous network conditions, RACS matches the performance of Multi-Paxos and Raft, achieving a throughput of 200k cmd/sec with a latency of 300ms, confirming that RACS introduces no unnecessary overhead. Finally, SADL-RACS —an optimized version of RACS designed for high performance and robustness—achieves an impressive throughput of 500k cmd/sec under synchronous network conditions and 196k cmd/sec under adversarial network conditions, further enhancing both performance and robustness.
Index Terms:
consensus, adversarial, networks.
I Introduction
Consensus [1] enables a set of replicas to maintain strongly consistent state while remaining resilient to failures, and is widely used in distributed cloud applications [2],[3],[4],[5],[6],[7],[8],[9],[10],[11],[12].
Most deployed consensus protocols in the cloud follow a leader-based design [3, 4, 13, 14]. The leader replica replicates client commands among the non-leader replicas and responds back to clients. Under normal case, where the network is synchronous, leader-based protocols provide moderate throughput and low latency.
Modern cloud networks [15, 16] use high-speed connections and are often synchronous. However, cloud networks can occasionally encounter adverse conditions such as high latency, packet loss, and partial connectivity [17]. These issues often stem from failing network hardware, congested links, inaccuracies in failure detection software, and flaws in monitoring systems [18, 19]. Hereafter, we refer to these occasional adverse cloud network conditions as ”adversarial network conditions”.
In practice, leader-based protocols and their multi-leader extensions [20][21][22][23] are designed to detect adversarial network conditions using timeouts [14]. If the leader replica does not respond within a pre-configured timeout, the remaining replicas elect a new leader replica. Electing a new leader replica often involves a separate ”view change” sub routine [14], where each replica votes on a potential next leader replica.
View change-based leader election introduces performance and robustness challenges [24, 25]. During view change, no new client commands are committed, resulting in a prolonged commit delay for commands submitted during this period. Client commands proposed during the view change accumulate into a substantial backlog, causing significant overhead that persists even after the completion of the view change-based leader election. Since each replica buffers client commands during the view change, a significant backlog of commands accumulates by the time the leader election is completed. The large backlog of buffered commands places significant strain on the next elected leader, leading to further performance degradation.
Asynchronous consensus protocols [25][26][27] ensure high robustness and performance under adversarial network conditions by leveraging randomization instead of relying on view change. However, to the best of our knowledge, asynchronous consensus protocols have never been deployed in cloud applications due to three main drawbacks: (1) they require significantly higher bandwidth compared to leader-based protocols [26]; (2) their design is often incompatible with the existing cloud software stack, hindering seamless integration and rapid adoption [27, 28]; and (3) their complexity makes them difficult to understand and implement, which limits industry adoption. Hence, most cloud providers continue to deploy leader-based consensus protocols despite their limited robustness against adversarial network conditions [2].
Problem Statement: Can we design a consensus protocol that (1) remains robust under adversarial network conditions, (2) achieves performance comparable to leader-based protocols in synchronous network conditions, and (3) maintains a simple design that allows easy integration with the existing cloud software stack?
This paper proposes RACS, a robust and easily understandable randomized consensus protocol that addresses the above problem. RACS achieves three key goals:
  • G1: Provides robustness against adversarial networks.
  • G2: Matches the performance of high speed leader-based protocols [13, 14] under synchronous network conditions.
  • G3: Employs an easy-to-understand design that seamlessly integrates with existing cloud software stack, enabling rapid adoption in the cloud.
Achieving robustness against adversarial network conditions has already been explored [26, 28, 25, 27], and we explicitly state that this is ”not” our primary contribution. Instead, RACS, our novel protocol, simultaneously achieves the three goals G1, G2, and G3—an accomplishment that no existing asynchronous protocol has achieved before. Designing RACS while achieving G1, G2 and G3 simultaneously is not trivial. First, designing a protocol that ensures robustness against adversarial network conditions without incurring additional overhead under synchronous conditions is a non-trivial algorithmic challenge. Second, designing a robust protocol while ensuring ease of understanding and compatibility with existing cloud software stacks is a challenge that no previous asynchronous consensus protocol has successfully addressed.
We implemented and evaluated RACS in Go [29] and compared it against the existing implementations of Multi-Paxos[13], Raft[14], and EPaxos [20]. We evaluated RACS on Amazon EC2 in a multi-region WAN setting. First, we show that RACS provides 28k cmd/sec of throughput, under adversarial network conditions, and outperforms Multi-Paxos and Raft which only provide 2.8k cmd/sec in the same setting. Second, we show that RACS delivers 200k cmd/sec in throughput under 300ms median latency, comparable to Multi-Paxos’s 200k cmd/sec throughput, under synchronous normal case conditions. Finally, we introduce novel optimizations for RACS and demonstrate that the optimized version; SADL-RACS achieves an impressive throughput of 500k cmd/sec.
This paper makes the following key contributions:
  • We propose RACS, the first practical randomized consensus protocol that simultaneously achieves (1) robustness against adversarial network conditions, (2) high performance under typical synchronous network conditions, and (3) seamless integration with existing cloud software stack.
  • We provide the formal proofs of RACS.
  • We propose SADL, an optimization for RACS that enables high throughput and high robustness.
  • We implement a prototype of RACS and conduct an experimental analysis on Amazon EC2.
II Background and Related Work
II-A Threat Model and Assumptions
[…]
We assume first-in-first-out (FIFO) perfect point-to-point links [1] between each pair of replicas. In practice, TCP[30] provides FIFO perfect point-to-point links. We assume a content-oblivious [31] network adversary; the adversary may manipulate network delays, but cannot observe the message content nor the internal replica state. In practice, TLS [32] encrypted channels between each pair of replica satisfy this assumption.
Due to the FLP impossibility result [33], any deterministic algorithm cannot solve consensus under asynchrony even under a single replica failure. In RACS, we circumvent the FLP impossibility result using randomization. In a typical cloud network, there are periods in which the network behaves synchronously, followed by phases where the network shows adversarial behavior. RACS makes use of this network behavior and operates in two modes: (1) synchronous mode and (2) fallback mode. During the synchronous periods, RACS employs a leader-based design to reach consensus using one round trip network delay, and during the fallback mode, RACS employs randomization.
II-B Consensus
Consensus enables a set of replicas to reach an agreement on a single history of values. A correct consensus algorithm satisfies four properties [1]: (1) validity: the agreed upon value should be previously proposed by a replica, (2) termination: every correct process eventually decides some value, (3) integrity: no process decides twice, and (4) agreement: no two correct processes decide differently.
Figure 1: High latency overhead due to view change in Raft.
Robustness problem in leader-based protocols: Multi-Paxos [13], Raft [14], and View Stamp Replication [34] employ a single leader replica to totally order commands. Under normal network conditions, the leader replica replicates client commands in one network round trip. When the network conditions become adversarial, leader-based protocols run a separate ”view-change” sub-routine that elects a new leader replica. View change introduces robustness challenges in leader-based protocols, which we illustrate using Figure 1.
In Figure 1, we run Raft—the most widely deployed consensus protocol in the cloud—with 5 replicas, with the view timeout configured at 5 seconds. At time t = 10s seconds, we simulate an adversarial network condition by introducing a 6-second delay on the links connected to the current leader. We first observe that the commands issued between $10$s and $15$s experience significantly high commit latency. This high latency arises from the view change, as no new commands are committed until the view change is fully completed. Second, we observe that commands issued after the completion of the view change also experience high commit latency. This is due to the backlog of commands that accumulated during the view change. Therefore, existing leader-based consensus protocols are unable to ensure robustness under adversarial network conditions. In contrast, RACS, our novel protocol, ensures robustness under adversarial network conditions.
Multi-leader protocols: Multi-leader protocols [21, 20, 23, 24] mitigate the leader bottleneck issue present in single-leader-based protocols. However, they still depend on view change subroutines to manage replica failures and network connectivity issues. Hence, under adversarial network conditions, multi-leader protocols also suffer from loss of performance. Other extensions of leader-based protocols [35][36][22][37] also lose liveness under adversarial network conditions due to their reliance on view change mechanisms. In contrast, our proposal RACS maintains liveness under adversarial network conditions.
Randomized consensus protocols: Randomized consensus protocols [38][25][27][26][39] overcome the adversarial network challenge by using randomization instead of view changes. However, they are rarely used in cloud applications due to (1) high performance overhead in synchronous network conditions, (2) incompatibility with the existing cloud software stack, and (3) complexity in understanding and implementing. In contrast, our protocol RACS provides a practical solution to adversarial network challenges by achieving: (1) robustness in adversarial network conditions, (2) optimal one round-trip commit latency in synchronous networks, and (3) a simple design for fast implementation and seamless industry adoption.
III RACS
III-A RACS Protocol Overview
Figure 2: Replicated log of RACS replicas.
Figure 3: Fallback mode of RACS.
We introduce RACS, a novel crash fault-tolerant randomized consensus algorithm. RACS achieves three important goals: (1) robustness against adversarial network conditions, (2) performance comparable to existing leader-based protocols under synchronous network conditions, and (3) simplicity in design, enabling rapid implementation and seamless integration into existing cloud software stack.
Given the wide adoption of Raft [14] in cloud software [40, 41], we design RACS by extending Raft. Building RACS on Raft as the foundation facilitates its rapid adoption in cloud environments, because existing cloud software stacks already support the programming abstractions and infrastructure required by Raft.
As illustrated in Figure 2, similar to Raft [14], each replica in RACS maintains a sequential log of blocks. Each block points to the previous block using a pointer. The purpose of RACS protocol is to ensure that each replica commits the same value for each block, despite replica and network failures.
Similar to Raft, RACS employs a leader-based design during the synchronous network conditions. The leader replica in RACS collects a batch of client commands, and replicates them in at least a majority of the replicas. Replication under synchronous network conditions takes one network round trip, the theoretical optimal. As long as the network conditions remain synchronous and the leader replica doesn’t crash, RACS commits commands in one network round trip.
Under adversarial network conditions, Raft employs a view change sub-routine, where each replica votes for a potential new leader replica. As shown in Figure 1, view change based leader election harms the performance. In RACS, we eliminate view change-based leader election, marking a fundamental difference from Raft. Instead of view-change based leader election, RACS employs a novel randomized fallback mode that keeps committing new client commands. Replacing view change sub routine with the randomized fallback mode of operation is the key difference between Raft and RACS.
Figure 3 illustrates the randomized fallback mode of RACS from the perspective of a single replica. In Figure 3, the network remains synchronous until block B3 is replicated. Due to an adversarial network condition, the leader replica fails to transmit B4, and therefor all the replicas ”automatically” enter the randomized fallback mode of operation. During the randomized fallback mode, each replica, independent from other replicas, act as a leader, and proposes new blocks. As shown in Figure 3, in a 3 replica configuration, the randomized fallback mode results in 3 parallel chains of blocks. The number of blocks proposed during the randomized fallback mode can be configured dynamically, but for the purpose of illustration, in this example we use two fallback blocks. Once each replica has proposed 2 fallback blocks, they consistently chose a single chain (the last fallback chain in Figure 3) out of all available fallback chains. If adversarial conditions persist, another round of fallback mode will be triggered. Otherwise, the synchronous path leader of the next view will replicate blocks in a single network round trip, extending the chosen fallback chain.
Replacing view change in Raft with a randomized fallback mode helps RACS achieve the three goals mentioned above. First, because new commands are committed in the randomized fallback mode, RACS remains robust under adversarial network conditions. Second, during the synchronous periods of execution, RACS uses a leader-based design, and incurs the same overhead as Raft. Finally, because RACS builds on Raft, it is easy to understand and requires fewer additions to the existing Raft implementations [42, 41], enabling rapid adoption in the cloud. RACS is the first protocol to simultaneously achieve all 3 goals.
III-B RACS Algorithmic Challenges
RACS has a simple design but faces non-trivial challenges in ensuring agreement and termination. First, depending on the network conditions, RACS must automatically switch between the synchronous single-chain mode and the fallback multi-chain mode, without manual intervention. Second, RACS must ensure that each replica commits the same chain of blocks, even when multiple chains exist during adversarial periods – each replica should consistently choose the same fallback chain at the end of the fallback mode. Section III-C addresses these challenges and presents the RACS algorithm.
Figure 4: RACS Algorithm.
III-C RACS Algorithm
Similar to Raft, RACS progresses as a sequence of views $v$ and rounds $r$ where each view has one or more rounds. A view consists of (1) a synchronous mode with a predefined leader $L_{\text{vcur}}$ and (2) a fallback mode of operation. In Figure 3 blocks $B1$$B5$ are from the same view, where blocks $B1$-$B3$ are synchronous mode blocks and blocks $B4$-$B5$ are fallback mode blocks. Blocks $B6$$B8$ are from the next view. A round represents the successive log positions in the replicated log. Round numbers are monotonically increasing. The pair $(v,r)$ is called a rank.
[…]
Fig. 4 depicts the pseudo-code of RACS. In the following discussion, when we say replica $p_{i}$ delivers a block $B$ from replica $p_{j}$, we imply that replica $p_{i}$ delivers $B$ and the causal history of $B$.
Synchronous mode: The synchronous mode of RACS is a leader-based and is similar to Raft, as depicted in lines 1–11 of Figure 4.. The synchronous mode leader $L_{vcur}$ for each view $v$ is predetermined and known to all replicas on bootstrap, for example $L_{vcur}=v_{cur}\%n$.
The synchronous mode begins either at the very start of the protocol or after a fallback mode has ended. Upon collecting a $n/2+1$ number of $<$vote$>$ messages, the leader replica forms a new block $B$, that extends the $block_{high}$ (see line 3 of Figure 4). The leader then broadcasts a $<$propose$>$ message for the block $B$ and the reference of the last committed block $block_{commit}$ (see line 4 of Figure 4).
Each replica $p_{i}$ delivers $<$propose, B, $block_{c}$$>$ message, if the rank of $B$ is greater than the rank of $p_{i}$ and if $p_{i}$ is in the synchronous mode of operation. If these two conditions are met, then $p_{i}$ first updates its $block_{high}$, $v_{cur}$, $r_{cur}$ (see line 6–7 of Figure 4) and then commits the block $block_{c}$ (see line 8 of Figure 4). $p_{i}$ then sends $p_{i}$’s $<$vote$>$ for $B$ to $L_{vcur}$ (see line 9 of Figure 4).
Upon receiving $n/2+1$ $<$vote$>$ messages for $B$ (see line 1 of Figure 4), the leader replica commits $B$ (and the causal history). As long as the synchronous mode leader $L_{vcur}$ is responsive, the synchronous mode will continue, committing a new block in one network round trip.
Each replica has a timeout clock which is reset whenever the replica receives a new $<$propose$>$ message (see line 5 of Figure 4). If the timeout expires (see line 11 of Figure 4), however, they will broadcast a $<$timeout$>$ message containing the $block_{high}$.
Fallback mode: The number of blocks proposed during the fallback mode is a tunable parameter that we explain in Section IV-A. For simplicity of illustration, we use 2 fallback blocks per view in the following description. Upon receiving $n/2+1$ $<$timeout$>$ messages, RACS enters the fallback mode of operation. In the fallback mode, all replicas act as leaders, concurrently. Each replica takes the highest $block_{high}$ they are aware of, forms a level $1$ fallback block $B$ with a monotonically increasing rank compared to the highest $block_{high}$ it received and sends a $<$propose-async$>$ message (see line 12–16 of Figure 4).
Upon receiving a $<$propose-async$>$ message from $p_{j}$, each replica $p_{i}$ sends back a $<$vote-async$>$ message to $p_{j}$ if the rank of the proposed level $1$ block is greater than the highest-ranked block witnessed so far (see line 17 of Figure 4).
Upon receiving $n/2+1$ $<$vote-async$>$ messages for the level $1$ fallback block $B_{f}$, each replica will send a level $2$ fallback block $B$ (see line 19 of Figure 4). The algorithm allows catching up to a higher ranked block by building upon another replica’s level $1$ block. This is meant to ensure termination for replicas that fall behind. All replicas, upon receiving a $<$propose-async$>$ message for a level 2 fallback block from $p_{j}$ send a $<$vote-async$>$ message to $p_{j}$ (see line 17–18 of Figure 4).
Once $n/2+1$ $<$vote-async$>$s have been gathered for the level $2$ fallback block $B_{f}$, each replica $p_{i}$ broadcasts an $<$asynchronous-complete$>$ message (see line 20 of Figure 4).
Once $n/2+1$ replicas have submitted $<$asynchronous-complete$>$ messages, the fallback mode is complete. The remaining task is to consistently select one chain from the set of fallback chains. This step is crucial for maintaining agreement. If two replicas select different fallback chains, the system loses the guarantee that all replicas commit the same sequence of blocks. To ensure that each replica consistently chooses the same chain, we employ a common-coin, an existing technique [26, 39, 25] used in distributed algorithms.
Common-coin: For each view $v$, common-coin($v$) returns a positive integer in the range (00, $n-1$) where $n$ is the total number of replicas. The invocation of common-coin($v$) for a given view $v$, at each replica returns the same value.
RACS uses a common-coin mechanism, where each replica runs a consistent random number generator with a pre-agreed seed. When used with the same seed, consistent random number generation protocols produce the same output sequence. Common-coin enables RACS to consistently choose a fallback chain after completing fallback mode.
Upon receiving $n/2+1$ $<$asynchronous-complete$>$ messages, each replica flips a common-coin (see line 21 of Figure 4) to get the elected leader Lelected. Each replica commits a level $2$ fallback block $B$ from Lelected if $B$ arrived amongst the first $n/2+1$ $<$asynchronous-complete$>$ messages (see line 22 of Figure 4). If a replica observes that the level $2$ fallback block from Lelected does not appear in the first $n/2+1$ $<$asynchronous-complete$>$ messages, but appears in Bfall[Lelected], then, the replica sets blockhigh to Bfall[Lelected] (see line 23 of Figure 4). If a replica observes that the level 2 fallback block from Belected is not in the first $n/2+1$ $<$asynchronous-complete$>$ messages and Bfall[Lelected] is null, then the replica doesn’t update its blockhigh and blockcommit. All replicas then exit the fallback mode and resume the synchronous path by uni-casting a $<$vote$>$ message to the synchronous leader Lvcur of the next view (see line 27 of Figure 4).
RACS protocol summary: Based on the above explanation of the RACS protocol, its outcomes can be summarized as follows. First, by committing new commands during the fallback mode, RACS ensures robustness against adversarial network conditions. Second, RACS commits new client commands in one network round trip, and hence provides optimal performance under synchronous normal case conditions. Finally, since RACS follows a simple design and extends Raft, it enables seamless integration with the existing cloud software stack.
III-D Correctness and Complexity
We now give the proof intuitions of RACS. The formal proofs have been deferred to the anonymous online appendix 111online appendix link due to page limit.
RACS only commits blocks proposed by replicas, thereby inherently satisfy the validity property. Because RACS does not commit the same block more than once, it trivially satisfies the integrity property. In the following sections, we focus on the agreement and termination properties.
Agreement property: RACS satisfies the agreement property by ensuring that if a block $B$ is committed in any round $r$, then all the blocks with round $r^{\prime}\geq r$ will extend $B$.
Proof: A block $B$ can be committed in 2 instances; (1) synchronous mode and (2) fallback mode.
In the synchronous mode, a block is committed by Lvcur when a majority of replicas vote for the synchronous block in the same round. Hence, if the leader commits a block $B$, then it is guaranteed that at least a $n/2+1$ majority of the replicas have $B$ as their $block_{high}$, or have $B$ in the causal history of $block_{high}$. Due to quorum intersection, all the future blocks will extend $B$ because each new block extends blockhigh.
In the fallback mode, the only fallback block $B$ of level $2$ sent by Lelected is committed if it is among the first $n/2+1$ $<$asynchronous-complete$>$ messages received in that view, which implies that a majority of the replicas have received and saved $B$ in Bfall[Lelected]. Hence if at least one node commits $B$ in the fallback mode, then it is guaranteed that at least a $n/2+1$ number of the replicas set $B$ as $block_{high}$, thus extending $B$ in the next round.
Therefore, we claim that agreement property holds in RACS.
Termination property: RACS ensures termination under both synchronous and adversarial network conditions.
Proof: During the synchronous mode, commands are committed in each round by the leader after receiving $<$votes$>$ from $n/2+1$ replicas, thereby satisfying the termination property.
In the fallback mode, if more than $n/2+1$ replicas enter the fallback mode, i.e. there are less than $n/2$ replicas on the synchronous mode, i.e. the synchronous path cannot progress anymore, therefore all replicas eventually enter the fallback mode. The fallback mode (with at least $n/2+1$ correct replicas in it) will eventually reach a point where $n/2+1$ correct replicas have sent $<$asynchronous-complete$>$ messages. The common-coin($v$) will therefore take place. If the result of the common-coin($v$) lands on one of the first $n/2+1$ replicas to have submitted an $<$asynchronous-complete$>$ message, then a new block is committed. Since the coin is unbiased, there is a probability $p$$>$$\frac{1}{2}$ that a new fallback block is committed in a given fallback mode of execution. Each replica therefore commits a new block in under 2 views in expectation and commits eventually with probability 1, thus ensuring the protocol’s termination.
Complexity The synchronous mode of RACS has a linear message and bit complexity for committing a block. The fallback mode of RACS has a complexity of $O(n^{2})$.
IV RACS Optimizations
IV-A Tuning fallback blocks
In Section III, we presented the RACS protocol using exactly two blocks in the fallback mode. In practice, adversarial network conditions can persist for unpredictable periods, making the use of a fixed number of blocks per view in fallback mode suboptimal.
In RACS, we dynamically adjust the number of blocks in fallback mode using Multi-Armed Bandit techniques [43] to optimize performance under varying network conditions. We initially run RACS with 2 blocks per fallback mode, and each replica monitors s: the number of times the algorithm switches between synchronous mode and the fallback mode. A high s implies a network where adversarial conditions persist, whereas a low s indicates a transient network problem that has self-healed. RACS employs a Multi-Armed Bandit-based explore/exploit strategy, increasing the number of blocks per fallback mode when s is high and decreasing it when s is low. RACS uses the same technique as Raft’s reconfiguration to consistently configure s across all replicas. Further details of this approach are beyond the scope of this paper and are omitted due to space constraints.
IV-B Application specific optimizations
Cloud applications that employ consensus protocols can be broadly categorized into two groups based on their performance objectives: (1) Class 1: applications that require moderate throughput and low latency [2] [44] [3] and (2) Class 2: applications that require high throughput and can tolerate moderately high latency [45] [46].
IV-B1 Optimization for class 1 applications
To reduce latency for class 1 applications, we employ pipelining, a well-established technique in distributed algorithms [25, 13, 20]. In pipelined-RACS, the synchronous mode leader $L_{\text{vcur}}$ proposes multiple blocks before receiving $n/2+1$ $<$vote$>$ messages. This allows multiple blocks to be committed within a single round-trip network delay during synchronous network periods, thereby reducing commit latency.
IV-B2 Optimization for class 2 applications
Figure 5: SADL Algorithm.
RACS blocks contain two types of data: (1) metadata about the block, including $block_{high}$, $block_{commit}$, $v$, and $r$, and (2) a batch of client commands. Typically, the metadata is only a few bytes in size, while the batch of client commands can span several kilobytes. Bundling both metadata and client commands within RACS blocks increases the block size, which limits the system’s throughput in synchronous mode due to the bandwidth bottleneck at the leader replica. This is a significant limitation of RACS, particularly for class 2 applications that demand high throughput.
To address this limitation, we propose SADL, a novel command dissemination layer that decouples command dissemination from the critical path of the consensus protocol. SADL is a consensus-agnostic command dissemination layer designed to reliably distribute client commands to at least a majority of replicas, ensuring efficient and fault-tolerant command propagation. Once SADL replicates commands across a majority of replicas, RACS layer no longer needs to include the commands in it’s blocks. Instead, RACS only embeds a small, fixed-size ”reference” to the batch already disseminated by SADL. Using a fixed-size reference instead of the entire command batch significantly reduces the RACS leader replica’s bandwidth consumption, leading to a substantial improvement in throughput, as demonstrated in Section VI.
Decoupling command dissemination from consensus has already been explored in the context of blockchain protocols [28, 47]. However, our SADL design significantly differs from existing approaches. Unlike existing decoupling mechanisms [28, 47], where the dissemination layer advances in a lock-step fashion, SADL allows each replica to build its own chain of SADL blocks independently of the pace of other replicas. This key difference enables SADL to achieve higher throughput than existing approaches. We present this subtle yet non-trivial distinction as a novel systems contribution of this paper.
SADL Protocol: Figure 5 depicts the SADL pseudo code. In SADL, each replica concurrently proposes SADL batches. For clarity of presentation, we describe the algorithm from the perspective of a single replica $p_{i}$. Upon receiving a batch of commands, replica $p_{i}$ creates a new SADL batch B, extending its last created SADL batch $B_{parent}$, and broadcasts a $<$SADL-Batch, B$>$ message (see lines 1–4 of Figure 5). Other replicas, upon receiving a new $<$SADL-Batch, B$>$, store B in chains, update tips to the round number of the parent of B, and send a $<$SADL-vote$>$ to $p_{i}$ (see lines 5–7 of Figure 5). Replica $p_{i}$, upon receiving $n/2+1$ $<$SADL-vote$>$ messages for its last proposed block B, updates the tips (see line 9 of Figure 5) and starts proposing a new SADL-batch (see lines 1–4 of Figure 5).
Replica $p_{i}$ broadcasts the next SADL batch only after receiving $n/2+1$ votes for the previous SADL batch it proposed. Hence, any replica, upon receiving a new SADL batch $B$ from $p_{i}$, can be certain that all SADL batches from $p_{i}$, prior to $B$ have already been replicated in at least a majority of the replicas. This property of SADL enables it to reference a large batch of client commands using a fixed-size integer array of several bytes. The tips integer array of size n stores the last successfully committed SADL batch for each replica, thereby allowing the entire set of client batches to be represented using a fixed-size integer array.
Using SADL with RACS: With SADL in place, RACS blocks no longer need to carry the heavy client batches and instead include only the tips integer array as the proposal for consensus. Upon committing a block in RACS, RACS commits all uncommitted batches in all SADL chains up to the last block pointed to by the tips array, effectively committing a consistent cut of the SADL chains. Since tips is an array of n elements, the size of RACS blocks is drastically reduced, effectively eliminating the leader bottleneck problem in the RACS synchronous mode. As we show in Section VI, SADL significantly increases throughput, albeit with a modest increase in latency. Therefore, for class 2 applications that require high throughput but can tolerate moderately high latency, SADL is well-suited.
V Implementation
We implemented pipelined-RACS and SADL-RACS using Go version 1.18 [29], in 3661 and 4631 lines of codes, respectively, as counted by CLOC [48]. We use the Go network library and TCP [30] for reliable FIFO point-to-point links between replicas. We used Protobuf encoding [49] for message serialization.
Both RACS and SADL-RACS implement batching in both clients and replicas as in existing implementations of Multi-Paxos, Raft[42], and EPaxos [50].
VI Experimental evaluation
This evaluation demonstrates the following 4 claims.
  • C1: RACS offers robustness against adversarial network conditions.
  • C2: Under synchronous network conditions, RACS performs comparably to existing leader-based algorithms.
  • C3: SADL-RACS delivers high throughput for class 2 applications that demand high throughput.
  • C4: SADL improves the scalability of RACS w.r.t (C4.1) increasing replica count and (C4.2): increasing payload size.
Since adversarial networks are much more common in the wide-area network (WAN) than in the local-area network (LAN), we focus on the WAN deployments in our evaluation, however, for completeness of experiments, we also compared the performance of RACS in a LAN in Section VI-D.
We compare RACS’s and SADL-RACS’s performance against three state-of-the-art SMR algorithms: Raft [14] [42], Multi-Paxos [13] [42], and EPaxos [20] [50]. Raft and Multi-Paxos are the most widely used consensus algorithms in cloud environments. EPaxos is a multi-leader protocol that enables parallel commits of non-interfering commands.
Setup: We test both a WAN setup where the replicas and clients are distributed globally across AWS cloud regions Sydney, Tokyo, Seoul, Osaka, and Singapore and a LAN setup where all replicas and clients are located in North California AWS cloud.
We use Amazon EC2 virtual machines [51] of type t2.xlarge (4 virtual CPUs, 16 GB memory) for replicas and clients, for WAN experiments. For LAN experiments, we use instances of type c4.4xlarge (16 virtual CPUs, 30 GB memory) for replicas and clients. We use Ubuntu Linux 20.04.5 LTS [52].
Workload and Benchmarks: Following the existing implementations Multi-Paxos and Raft [42], we use a $map[string]string$ key-value store and Redis[53] as backend applications.
In our experiments, we use $n$ replicas and $n$ clients. Clients generate client commands with a Poisson distribution in the open-loop model [54]. All algorithms employ batching in both clients and replicas. A single client command is a 17 bytes string: 1-byte GET/PUT opcode plus 8-byte keys and values, consistent with command sizes used in prior research and production systems [39, 11].
For RACS, SADL-RACS, Multi-Paxos, and Raft we measure the client observed end-to-end execution latency, which accounts for the latency overhead for total ordering and executing commands. EPaxos provides two modes of operations: (1) partial ordering of commands without execution (denoted “EPaxos-commit” in the graphs) and (2) partial ordering of commands with execution (denoted “EPaxos-exec” in the graphs). Trivially, “EPaxos-commit” outperforms RACS and SADL-RACS because ”EPaxos-commit” only provides a partial order of commands, which enables higher parallelism. Hence, ”EPaxos-commit” provides an apples-to-oranges comparison, however, we present the results in this evaluation, for completeness. We also found and reported bugs in the existing implementation of EPaxos code that prevent execution under adversarial network conditions, crashes, and when deployed with more than 5 replicas. Hence we use EPaxos only under normal-case performance evaluation.
We measure throughput in commands per second (cmd/sec), where a command is one 17-byte command. We measure the latency in milliseconds.
Figure 6: Adversarial Performance in the WAN with 5 replicas – note that Multi-Paxos and Raft lines overlap
VI-A Asynchronous Performance
This experiment evaluates RACS and SADL-RACS under simulated network attack scenarios, similar to attacks used in  [24, 25]. Our simulated attacker increases the egress packet latency of a minority of replicas chosen at random, by 500ms, dynamically, in time epochs. This experiment runs in the WAN setting with 5 replicas and 5 clients. We depict the results in Figure 6.
RACS vs {Multi-Paxos and Raft}:We observe that RACS provides 28k cmd/sec saturation throughput, in contrast, Multi-Paxos and Raft have saturation throughput at 2.8k cmd/sec. Under adversarial network conditions Multi-Paxos and Raft undergo repeated view changes, and fail at successfully committing commands. In contrast, due to the fallback mode of execution, RACS provides liveness under adversarial network conditions. Hence we prove the claim C1.
RACS vs SADL-RACS: We observe that SADL-RACS delivers 196k saturation throughput, thus providing 168k cmd/sec more throughput than RACS. Because SADL-RACS eliminates the leader bottleneck and the SADL protocol is robust against adversarial network conditions (as it does not rely on timeouts), SADL-RACS achieves higher throughput than RACS, even under adversarial network conditions.
(a) Median Latency 3 replicas
(b) Median Latency 5 replicas
(c) 99% Latency 5 replicas
(d) Median Latency 11 replicas
Figure 7: Throughput versus latency for WAN normal-case execution, comparing pipelined RACS and SADL-RACS to pipelined Multi-Paxos, and pipelined EPaxos, with 3, 5 and 11 replica ensembles.
VI-B RACS WAN Normal Case Performance
In this experiment, we evaluate the normal-case synchronous performance of RACS deployed in 5 geographically distant AWS regions. Because we observed the same throughput and latency for Raft and Multi-Paxos (due to their identical structure during normal-case replication), we only depict the Multi-Paxos performance for clarity in the figures. Fig. 7(b) and Fig. 7(c) depict the experimental results using 5 replicas and 5 clients.
RACS vs Multi-Paxos: We observe in Fig. 7(b) that RACS delivers a saturation throughput of 200k cmd/sec throughput under 300ms median latency, which is comparable to the performance of Multi-Paxos (200k cmd/sec under 300ms latency). In the synchronous execution both RACS and Multi-Paxos have 1 round trip latency per batch of commands, hence share the same performance characteristics. Hence the experimental claim C2 holds.
RACS vs Epaxos-commit: We observe in  Fig. 7(b) that EPaxos-commit (without command execution) delivers a throughput of 500k+ cmd/sec under 170ms median latency. The EPaxos-commit experiment employs a conflict rate of 2%[55] hence 98% of the time, commands are committed in one round trip, without serializing through a leader replica. In contrast, RACS builds a total order of commands, serialized using a leader-replica, hence naturally the RACS performance is bottlenecked by the leader replica’s capacity.
RACS vs Epaxos-exec: As shown in Fig. 7(b), the median latency of EPaxos-exec (with command execution) is 300ms higher on average than RACS in the 50k–200k cmd/sec throughput range. This higher latency stems from EPaxos’s dependency management cost [55, 56]. Hence, we conclude that when measured for execution latency, RACS outperforms EPaxos.
RACS vs SADL-RACS: As shown in Fig. 7(b), SADL-RACS achieves a throughput of 500k cmd/sec, exceeding RACS by 300k cmd/sec, however, with an additional latency cost of 500ms. These results demonstrate the effectiveness of SADL in eliminating the leader bottleneck in RACS for class 2 applications that require high throughput but can tolerate moderately high latency, supporting the claim C3.
(a) Median Latency (17B)
(b) Median Latency (73B)
(c) Median Latency (265B)
(d) 99% Latency (73B)
Figure 8: Throughput versus latency for WAN normal-case execution, comparing SADL-RACS to pipelined Multi-Paxos and pipelined RACS using 17B, 73B and 265B command sizes, using 5 replicas
VI-C Scalability of SADL
In this experiment, we aim to quantify the scalability of SADL-RACS. We consider two factors of scalability: (1) scalability w.r.t increasing replication factor and (2) scalability w.r.t increasing payload size.
Although we conducted experiments with both Multi-Paxos and Raft, we observed the same throughput and latency behavior due to their identical protocol structure during normal-case replication (both replicate in a single network round trip, serialized through a leader). Therefore, for clarity in figures and explanation, we present only the Multi-Paxos results.
Scalability w.r.t increasing replication factor: In this experiment, we evaluate the scalability of SADL-RACS by running it with an ensemble of three (minimum replication allowed), five (common replication factor) and eleven replicas (improved robustness to concurrent replica failures), located in geographically separated AWS regions. Fig. 7 compare the scalability of SADL-RACS against pipelined RACS, pipelined Multi-Paxos, and pipelined EPaxos, for different replication factors.
RACS vs {Multi-Paxos and Raft}: We observe that the saturation throughput of Multi-Paxos and RACS decreases from 230k to 130k cmd/sec (under 600ms median latency) when the replication factor is increased from 3–11. With increasing replica count, the leader replica in RACS and Multi-Paxos has to send and receive more messages, due to increased quorum sizes, hence the performance is bottlenecked by the leader’s bandwidth capacity.
RACS vs SADL-RACS: We observe that SADL-RACS provides a throughput of 380k cmd/sec (under 600 ms median latency), when the replication factor is 11. SADL-RACS outperforms pipelined-RACS and pipelined Multi-Paxos by 192% in the 11 replica scenario. This confirms that SADL enables RACS to increase throughput by avoiding the leader bottleneck through the decoupling of command dissemination from consensus. Hence we prove the claim C4.1.
Scalability w.r.t increasing payload size: In this experiment, we evaluate the impact of payload size for the SADL-RACS performance. We experiment with 3 key sizes: 8B, 64B, and 256B, used in recent SMR work [57]. Combined with 1B opcode and 8B value, these key sizes result in 17B, 73B, and 265B command sizes. We deploy SADL-RACS, pipelined Paxos and pipelined-RACS in a WAN setting with 5 replicas and 5 clients. Fig. 8 depicts the results.
We observe that for each command size, the saturation throughput of SADL-RACS is at least 2 times the throughput of RACS and Multi-Paxos. With increasing command size, the leader replica’s bandwidth of RACS (and Multi-Paxos) becomes the bottleneck. In contrast, thanks to the decoupling of command dissemination from consensus, SADL-RACS evenly distributes the bandwidth overhead among all the replicas, and sustains higher throughput. This proves the claim C4.2.
VI-D RACS LAN Normal Case Performance
Figure 9: Throughput versus latency for LAN normal-case execution, comparing RACS to Multi-Paxos, EPaxos, and Raft using 5 replicas
In this experiment, we quantify the performance of the RACS in a LAN setting. We conduct experiments with 5 replicas and 5 clients deployed in the same AWS region. Figure 9 illustrates the LAN performance of RACS.
RACS vs {Multi-Paxos and Raft}: We first observe that RACS achieves a saturation throughput of 420k cmd/sec, under a median latency upper bound of 10ms, that is comparable to the saturation throughput of Multi-Paxos (450k) and Raft (440k). Under normal case executions, RACS, Multi-Paxos, and Raft have 1 round-trip latency, serialized through a leader, hence provide comparable performance.
RACS vs EPaxos: Second, we observe that Epaxos-commit outperforms RACS, both in terms of latency and throughput. The EPaxos-commit experiment employs a conflict rate of 2%[55] hence 98% of the time, commands are committed in one round trip, without serializing through a leader replica. In contrast, RACS builds a total order of commands, serialized using a leader-replica, hence naturally the RACS performance is bottlenecked by the leader replica’s capacity.
VII Conclusion
We presented RACS, a novel randomized consensus protocol that simultaneously achieves three key goals. First, RACS ensures robustness under adversarial network conditions by utilizing a fallback mode of operation, replacing the traditional view change used by existing leader-based protocols. Second, RACS achieves performance comparable to existing leader-based protocols, under normal-case synchronous network conditions. Third, RACS adopts a simple design inspired by Raft, enabling easy engineering and seamless integration in cloud environments. RACS is the first protocol to simultaneously achieve these three goals.
We experimentally demonstrated RACS’s robustness and performance guarantees. RACS is suitable for adoption in the industry due to its robustness, high performance, and simplicity.
Appendix A RACS Formal Proofs
A-A Definition
elected-fallback block: We refer to an fallback block $B_{f}$ generated in view $v$ with level $2$ as an elected-fallback block, if the common-coin-flip($v$) returns the index of the proposer $p_{l}$ who generated $B_{f}$ in the view $v$ and if the $<$asynchronous-complete$>$ for $B_{f}$ exists in the first $n-f$ $<$asynchronous-complete$>$ messages received. An elected-fallback block is committed same as a synchronously-committed block.
A-B Proof of agreement
Theorem 1.
Let $B$ and $\tilde{B}$ be two blocks with rank $(v,r)$. Each of $B$ and $\tilde{B}$ can be of type: (1) synchronous block which collects at least $n-f$ votes or (2) elected-fallback block or (3) level $1$ fallback block which is a parent of an elected-fallback block. Then $\tilde{B}$ and $B$ are the same.
Proof.
This holds directly from the block formation – if both $B$ and $\tilde{B}$ has the same rank, then due to quorum intersection, there exists at least one node who voted for both blocks in the same rank, which is a contradiction to our assumption of non malicious nodes. ∎
Theorem 2.
Let $B$ and $\tilde{B}$ be two adjacent blocks, then $\tilde{B}.r=B.r+1$ and $\tilde{B}.v\geq B.v$.
Proof.
According to the algorithm, there are three instances where a new block is created.
  • Case 1: when $isAsync$ = false and $L_{vcur}$ creates a new synchronous block by extending the $block_{high}$ with rank ($v$,$r$). In this case, $L_{vcur}$ creates a new block with round $r+1$. Hence the adjacent blocks have monotonically increasing round numbers.
  • Case 2: when $isAsync$ = true and upon collecting $n-f$ $<$timeout$>$ messages in view $v$. In this case, the replica selects the $block_{high}$ with the highest rank $(v,r)$, and extends it by proposing a level $1$ fallback block with round $r+1$. Hence the adjacent blocks have monotonically increasing round numbers.
  • Case 3: when $isAsync$ = true and upon collecting $n-f$ $<$vote-async$>$ messages for a level $1$ fallback block. In this case, the replica extends the level $1$ block by proposing a level $2$ block with round $r+1$. Hence the adjacent blocks have monotonically increasing round numbers.
The view numbers are non decreasing according to the algorithm. Hence Theorem 2 holds. ∎
Theorem 3.
If a synchronous block $B_{c}$ with rank $(v,r)$ is committed, then all future blocks in view $v$ will extend $B_{c}$.
Proof.
We prove this by contradiction.
Assume there is a committed block $B_{c}$ with $B_{c}$.$r$ = $r_{c}$ (hence all the blocks in the path from the genesis block to $B_{c}$ are committed). Let block $B_{s}$ with $B_{s}$.$r$ = $r_{s}$ be the round $r_{s}$ block such that $B_{s}$ conflicts with $B_{c}$ ($B_{s}$ does not extend $B_{c}$). Without loss of generality, assume that $r_{c}<r_{s}$.
Let block $B_{f}$ with $B_{f}$.$r$ = $r_{f}$ be the first valid block formed in a round $r_{f}$ such that $r_{s}\geq r_{f}>r_{c}$ and $B_{f}$ is the first block from the path from genesis block to $B_{s}$ that conflicts with $B_{c}$; for instance $B_{f}$ could be $B_{s}$. $L_{vcur}$ forms $B_{f}$ by extending its $block_{high}$. Due to the minimality of $B_{f}$ ($B_{f}$ is the first block that conflicts with $B_{c}$), $block_{high}$ contain either $B_{c}$ or a block that extends $B_{c}$. Since $block_{high}$ extends $B_{c}$, $B_{f}$ extends $B_{c}$, thus we reach a contradiction. Hence no such $B_{f}$ exists. Hence all the blocks created after $B_{c}$ in the view $v$ extend $B_{c}$. ∎
Theorem 4.
If a synchronous block $B$ with rank $(v,r)$ is committed, an elected-fallback block $\tilde{B}$ of the same view $v$ will extend that block.
Proof.
We prove this by contradiction. Assume that a synchronous block $B$ is committed in view $v$ and an elected-fallback block $\tilde{B}$ does not extend $B$. Then, the parent level $1$ block of $\tilde{B}$, $\tilde{B_{p}}$, also does not extend $B$.
To form the level $1$ $\tilde{B_{p}}$, the replica collects $n-f$ $<$timeout$>$ messages, each of them containing the $block_{high}$. If $B$ is committed, by theorem 3, at least $n-f$ replicas should have set (and possibly sent) $B$ or a block extending $B$ as the $block_{high}$. Hence by intersection of the quorums $\tilde{B_{p}}$ extends $B$, thus we reach a contradiction. ∎
Theorem 5.
At most one level 2 fallback block from one proposer can be committed in a given view change.
Proof.
Assume by way of contradiction that 2 level $2$ fallback blocks from two different proposers are committed in the same view. A level $2$ fallback block $B$ is committed in the fallback phase if the common-coin-flip($v$) returns the proposer of $B$ as the elected proposer. Since the common-coin-flip($v$) outputs the same elected proposer across different replicas, this is a contradiction. Thus all level $2$ fallback blocks committed during the same view are from the same proposer.
Assume now that the same proposer proposed two different level $2$ fallback blocks. Since no replica can equivocate, this is absurd.
Thus at most one level $2$ fallback block from one proposer can be committed in a given view change. ∎
Theorem 6.
Let $B$ be a level $2$ elected-fallback block that is committed, then all blocks proposed in the subsequent rounds extend $B$.
Proof.
We prove this by contradiction. Assume that level two elected-fallback block $B$ is committed with rank $(v,r)$ and block $\tilde{B}$ with rank ($\tilde{v}$, $\tilde{r}$) such that ($\tilde{v}$, $\tilde{r}$) $>$ $(v,r)$ is the first block in the chain starting from $B$ that does not extend $B$. $\tilde{B}$ can be formed in two occurrences: (1) $\tilde{B}$ is a synchronous block in the view $v+1$ or (2) $\tilde{B}$ is a level $1$ fallback block with a view strictly greater than $v$. (we do not consider the case where $\tilde{B}$ is a level 2 elected-fallback block, because this directly follows from 1)
If $B$ is committed, then from the algorithm construction it is clear that a majority of the replicas will set $B$ as $block_{high}$. This is because, to send a $<$asynchronous-complete$>$ message with $B$, a replica should collect at least $n-f$ $<$vote-async$>$ messages. Hence, its guaranteed that if $\tilde{B}$ is formed in view $v$+1 as a synchronous block, then it will observe $B$ as the $block_{high}$, thus we reach a contradiction.
In the second case, if $\tilde{B}$ is formed in a subsequent view, then it is guaranteed that the level $1$ block will extend $B$ by gathering from the $<$timeout$>$ messages $B$ as $block_{high}$ or a block extending $B$ as the $block_{high}$, hence we reach a contradiction. ∎
Theorem 7.
There exists a single history of committed blocks.
Proof.
Assume by way of contradiction there are two different histories $H_{1}$ and $H_{2}$ of committed blocks. Then there is at least one block from $H_{1}$ that does not extend at least one block from $H_{2}$. This is a contradiction with theorems 34 and 6. Hence there exists a single chain of committed blocks. ∎
Theorem 8.
For each committed replicated log position $r$, all replicas contain the same block.
Proof.
By theorem 2, the committed chain will have incrementally increasing round numbers. Hence for each round number (log position), there is a single committed entry, and by theorem 1, this entry is unique. This completes the proof. ∎
A-C Proof of termination
Theorem 9.
If at least $n-f$ replicas enter the fallback phase of view $v$ by setting $isAsync$ to true, then eventually they all exit the fallback phase and set $isAsync$ to false.
Proof.
If $n-f$ replicas enter the fallback path, then eventually all replicas (except for failed replicas) will enter the fallback path as there are less than $n-f$ replicas left on the synchronous path due to quorum intersection, so no progress can be made on the synchronous path and all replicas will timeout. As a result, at least $n-f$ correct replicas will broadcast their $<$timeout$>$ message and all replicas will enter the fallback path.
Upon entering the fallback path, each replica creates a fallback block with level $1$ and broadcasts it. Since we use perfect point-to-point links, eventually all the level $1$ blocks sent by the $n-f$ correct replicas will be received by each replica in the fallback path. At least $n-f$ correct replicas will send them $<$vote-async$>$ messages if the rank of the level $1$ block is greater than the rank of the replica. To ensure liveness for the replicas that have a lower rank, the algorithm allows catching up, so that nodes will adopt whichever level $1$ block which received $n-f$ $<$vote-async$>$ arrives first. Upon receiving the first level $1$ block with $n-f$ $<$vote-async$>$ messages, each replica will send a level $2$ fallback block, which will be eventually received by all the replicas in the fallback path. Since the level $2$ block proposed by any block passes the rank test for receiving a $<$vote-async$>$, eventually at least $n-f$ level $2$ blocks get $n-f$ $<$vote-async$>$. Hence, eventually at least $n-f$ replicas send the $<$asynchronous-complete$>$ message, and exit the fallback path. ∎
Theorem 10.
With probability $p>\frac{1}{2}$, at least one replica commits an elected-fallback block after exiting the fallback path.
Proof.
Let leader $L_{elected}$ be the output of the common-coin-flip($v$). A replica commits a block during the fallback mode if the $<$asynchronous-complete$>$ message from $L_{elected}$ is among the first $n-f$ $<$asynchronous-complete$>$ messages received during the fallback mode, which happens with probability at least greater than $\frac{1}{2}$. Hence with probability no less than $\frac{1}{2}$, each replica commits a chain in a given fallback phase. ∎
Theorem 11.
A majority of replicas keep committing new blocks with high probability.
Proof.
We first prove this theorem for the basic case where all replicas start the protocol with $v=0$. If at least $n-f$ replicas eventually enter the fallback path, by theorem 9, they eventually all exit the fallback path, and a new block is committed by at least one replica with probability no less than $\frac{1}{2}$. According to the asynchronous-complete step, all nodes who enter the fallback path enter view $v=1$ after exiting the fallback path. If at least $n-f$ replicas never set $isAsync$ to true, this implies that the sequence of blocks produced in view $1$ is infinite. By Theorem 2, the blocks have consecutive round numbers, and thus a majority replicas keep committing new blocks.
Now assume the theorem 11 is true for view $v=0,...,k-1$. Consider the case where at least $n-f$ replicas enter the view $v=k$. By the same argument for the $v=0$ base case, $n-f$ replicas either all enter the fallback path commits a new block with $\frac{1}{2}$ probability, or keeps committing new blocks in view $k$. Therefore, by induction, a majority replicas keep committing new blocks with high probability. ∎
Theorem 12.
Each client command is eventually committed.
Proof.
If each replica repeatedly keeps proposing the client commands until they become committed, then eventually each client command gets committed according to theorem 11. ∎